# Evaluating Transparency of Machine Generated Fact Checking Explanations

### v1

@@ Appendix B Transparency Evaluation Details
Given annotators’ prediction $X_{pred}^{mask}$ and the reference label set $X_{cit}^{mask}$, $Precision$ reflects the proportion of reference label in annotator prediction:
@@
$F1$ combines $Precision$ and $Recall$ via harmonic mean.
@@
Entropy (Shannon, 1948) is usually introduced to measure the randomness and the degree of uncertainty in the system:
$H=-\sum{{p_{k}}\;{log}\;{p_{k}}}$ (4)
Appendix C Generation Statistics
Setting Model Claim Length Evidence Size Explanation Length
Human GPT4 22.32$\pm$10.89 3.16$\pm$1.50 140.58$\pm$39.18
GPT35
LLaMA2-70b
Machine GPT4 22.32$\pm$10.89 5.07$\pm$2.42 171.26$\pm$51.45
GPT35 4.68$\pm$2.58 175.89$\pm$56.99
LLaMA2-70b 6.24$\pm$3.39 214.02$\pm$98.27
Table 4: Statistics of the data and generated explanations. The “Setting” column indicates whether evidence is human- or machine-selected. Claim Length and Explanation Length refer to token length, which is tokenized by OpenAI’s tiktoken v0.7.0. Evidence Size refers to the number of selected evidence passages.
Table 4 shows generation statistics. Claim Length is the same for all settings since we sampled same 100 claims for all experiments. Machine tends to extract more evidences than for Human and it also generates longer explanations.
Appendix D Human Annotation Details
@@
The annotation contains two tasks. The first task provides annotators with selected evidence passages asks to find masked citation sentence in the explanation. The second task requires annotators to judge the utility of the explanation in clarifying the claim (Section 3).
@@
The annotation task was performed on Amazon Mechanical Turk (AMT)  [^9] https://www.mturk.com/. We applied pre-screening pilot studies to find qualified annotators. We conducted individual reviews of submitted annotation results and offered feedback to annotators to address any misconceptions or confusion about the task. Annotators who performed well in the pilot study were selected to participate in final human evaluation. In order to maintain high quality throughout the annotation, we deploy quality control to identify and remove extremely poor-performing annotators. We also maintained continuous communication with annotators to address questions. In total, 68 annotators participated in the final human evaluation.
@@
Quality control was implemented for both pilot study and the main task. In pilot study, each Human Intelligence Task (HIT) containing 3/6 control questions. 2 of them are positive questions containing exact one answer each and 1 is negative question containing no correct answer (the original answer sentence has been removed). Annotators are expected to choose “There isn’t any sentence that can correctly cite the highlighted core evidence.” in such cases. All control questions are manually inspected to avoid any confusion. An annotator who fails on any control questions will be disqualified from participating further tasks. In total, 9 pilot studies were released to recruit qualified annotators.
During the final human evaluation, as annotators had already participated in the pilot study, a different batch of control questions were deployed to mitigate potential bias. We used ratio of failing control questions as an indicator to identify and remove extremely poor-performing workers.
@@
Annotators were compensated \$1.81 (US Dollar) per claim, and \$0.38 (US Dollar) per query response pair for responses without citations. On average, annotators took approximately 10-12 minutes to complete two tasks.
@@
During human evaluation, each claim was annotated by 5 human workers. Since common metrics like Cohen’s kappa (McHugh, 2012) and Fleiss’s kappa are not applicable to our multi-label annotation scenario, we used Measuring Agreement on Set-valued Items distance (MASI) (Passonneau, 2006) together with Krippendorff’s alpha (Krippendorff, 2011) to compute annotator agreement. Finally, we achieved alpha value of 0.48 for GPT-4, 0.32 for GPT-35 and 0.33 for LLaMA2-70b instances.
Appendix E Prompts
Inspired by Gao et al. (2023), we developed and optimized prompt for our task. We show evidence selection prompt in Table 6 and explanation generation prompt in Table 5.
Prompt
Instructions: You are required to write an accurate, coherent and logically consistent explanation for the claim based on the given veracity and list of reasons in one paragraph. Use an unbiased and journalistic tone. When citing several search results, use [1][2][3]. Ensure that each reason is cited only once. Do not cite multiple reasons in a single sentence.
Reasons:
Reason [1] What’s more, the picture referenced in the Facebook post alleging that anglerfish are typically 7 feet is taken from the Australian Museum’s 2012 exhibit titled "Deep Oceans. Reason [2] When the exhibit opened in June 2012, The Sydney Morning Herald reported on how the exhibit’s team had created an "oversized anglerfish" and listed the many steps in making it: "Pieces such as the oversized anglerfish, with huge fangs and antenna-like flashing rod to attract prey, begin with cutting and welding a metal frame, then sculpting material over it and, finally, hand painting it," the story says.
Claim:The typical anglerfish is seven feet long.
Veracity: False
Explanation:
Table 5: Prompts for explanation generation
Appendix F Utility Score Calibration
@@
Appendix G Experiment Details
We ran all offline models on 4 Nvidia A100 GPUs in a data parallel fashion. Explanation generation with LLaMA2-7B takes around 2 hours and LLaMA2-70B takes around 4 hours. The cost of generating explanations was \$49 (US Dollar) with GPT-4 and \$10 (US Dollar) with GPT-35.
Appendix H Scientific Artifacts
@@
Table 6: Prompts for model evidence selection
Appendix I Annotation Interface
Figure 2: Annotation interface Page 1

### v2

@@ 3.3 Human Assessment Protocol
Given an annotator prediction $X_{pred}^{mask}$ and the reference label set $X_{cit}^{mask}$, $Precision$ reflects the proportion of reference label in annotator prediction:
@@
$F1$ combines $Precision$ and $Recall$ via the harmonic mean.
@@
3.4 Human Annotation Details
@@
We provide the annotator with the following information: a claim, veracity label, full evidence set, a selected evidence passage, and an explanation, and ask them to select (= highlight) sentences in the explanation that should cite the selected evidence passage (as in Figure 2). See Appendix F for the full annotation guidelines and interface. [^5] We also ask the annotators to rate the utility of the explanation after the citation recovery task, but we did not use these ratings for the experiments in this paper. Note that for each claim, we only select one random evidence passage to do citation recovery. For example in Figure 2, we select citation 9 to mask in the explanation. We address this limitation when we introduce LLM as annotator in Section 4, where every evidence passage is evaluated.
@@
The annotation task was done on Amazon Mechanical Turk (AMT). [^6] https://www.mturk.com/ We applied pre-screening pilot studies to find qualified annotators. We conducted individual reviews of submitted annotations, and offered feedback to annotators to address any misconceptions or confusion about the task. Annotators who performed well in the pilot study were invited to do a final round of annotation that produced the annotated data. In order to maintain high quality throughout the annotation, we used quality control questions to identify and remove poor-performing annotators. In total, we recruited 68 annotators for the final round of annotation.
@@
Quality control was implemented for both the pilot study and final study. In the pilot study, each Human Intelligence Task (HIT) contained 3/6 control questions, 2 of which were positive questions containing exactly one answer each, and 1 of which is a negative question containing no correct answer (the original answer sentence has been removed). Annotators are expected to choose “There isn’t any sentence that can correctly cite the highlighted core evidence.” in such cases. All control questions were manually validated by the first author of the paper. An annotator who fails on any control questions is disqualified from participating in further tasks. In total, 9 pilot studies were released to recruit qualified annotators. For the final study, we use a different batch of control questions (i.e. there is no overlap in control questions between the pilot and final study). Workers with an accuracy of below 70% were removed and prevented from doing more tasks.
@@
Annotators were paid USD\$2.70 for each HIT. During the pilot study, workers were paid the base rate of \$1 and a bonus of \$1.70 if they passed the quality control. On average, a HIT took approximately 7 minutes to complete, resulting in a salary rate of \$15 per hour. In the final study, all workers were compensated for their time, irrespective of whether they passed the quality control.
Figure 3: The inner-annotator agreement of human annotation. “Evi Src” indicates the whether the evidence is human or machine selected.
@@
Each claim was annotated by 5 annotators. As agreement metrics like kappa (McHugh, 2012) are not applicable to our multi-label scenario, we use Krippendorff’s alpha (Krippendorff, 2011) to compute annotator agreement. [^7] For Krippendorff’s alpha, we use Jaccard distance to measure distance between two annotations (recall that each annotation result is a set, and as such we use Jaccard distance to measure the degree of set overlap).
Figure 3 presents the inter-annotator agreement results. On average, GPT-4 explanations produce the best agreement, followed by GPT-3.5. LLaMA2 has the lowest agreement, because its generated explanations tend to be quite noisy (we will revisit this in the next section). Interestingly, explanations generated using human-selected evidence have a lower agreement score. Overall, these agreement results indicate moderate to good agreement for GPT-4 and GPT-3.5.
3.5 Human Evaluation Results
Given the human annotation, we now compute citation recovery performance (set precision, recall, and F1 in Section 3.3) to assess attribution quality in the generated explanations. Results are presented in Table 3.
GPT-4 has the best performance by a comfortable margin, suggesting that its generated explanations cite their sources more accurately than the other LLMs. Manual analysis reveals that LLaMA2 explanations often contain incorrect citations, such as [1][2][3] which are directly copied from the prompt/instruction, as well as random links. That said, even the best model (GPT-4 with machine-selected evidence) only produces an F1 of approximately 0.74, which means a good proportion of citations are still not faithful. There’s also substantial variance ($\pm$0.31), indicating the performance is far worse in the worst case scenario.
Evi Src Gen Precision Recall F1
Human GPT-4 0.62$\pm$0.29 0.67$\pm$0.29 0.63$\pm$0.29
GPT-3.5 0.52$\pm$0.29 0.59$\pm$0.30 0.52$\pm$0.29
LLaMA2 0.48$\pm$0.31 0.52$\pm$0.32 0.49$\pm$0.31
Machine GPT-4 0.72$\pm$0.32 0.79$\pm$0.31 0.74$\pm$0.31
GPT-3.5 0.55$\pm$0.39 0.55$\pm$0.39 0.53$\pm$0.37
LLaMA2 0.49$\pm$0.39 0.51$\pm$0.40 0.49$\pm$0.38
Table 3: Human evaluation results of generated explanations. The “Evi Src” column indicates whether evidence is human- or machine-selected.
Looking at human-selected vs. machine-selected evidence for a given LLM, we generally see similar attribution performance, implying that we may not need perfectly curated evidence for generating explanations. To understand how different the machine-selected evidence is to the human-selected evidence, we compute set precision, recall and F1 using the human-selected evidence as ground truth, and present the results in Table 4. Interestingly, across all LLMs we see only see a low to moderate F1 (0.3–0.5), meaning the machine-selected evidence is indeed quite different to human-selected evidence, but despite this difference, the resulting generated explanations do not differ much in terms of attribution quality.
Retriever Precision Recall F1
GPT-4 0.40$\pm$0.21 0.75$\pm$0.23 0.47$\pm$0.20
GPT-3.5 0.34$\pm$0.22 0.60$\pm$0.35 0.39$\pm$0.22
LLaMA2-70b 0.29$\pm$0.18 0.68$\pm$0.32 0.36$\pm$0.18
Table 4: Evidence retrieval performance.
4 Automatic Evaluation of Evidence Attribution
We now attempt to automate the annotation process, i.e. automate the citation recovery task in the assessment protocol (Section 3.3). We experiment with two approaches: Natural Language Inference (NLI), and LLM annotators.
NLI
Most existing work frames the evaluation of citation quality as an NLI task (Gao et al., 2023). That is, given a pair of (premise, hypothesis), the NLI model outputs a label indicating whether the premise entails or contradicts the hypothesis. In our scenario, the premise is an evidence passage ($e_{k}$) while the hypothesis is an explanation sentence (a sentence in $X$). Given a selected evidence passage, we pair it with every sentence in the explanation and use a pretrained NLI model to detect instances of entailment: any sentence that is predicted to be entailed by the evidence passage is added to the recovered set ($X_{pred}^{mask}$).
In terms of pretrained NLI models, we use two models: DeBERTaV3 (He et al., 2023) and TRUE (Honovich et al., 2022). [^8] Detailed model versions: DeBERTaV3 = deberta-v3-base, TRUE = t5_xxl_true_nli_mixture. The output label set for DeBERTaV3 is “entailment”, “neutral” and “contradiction”, and for TRUE is “1” (entailment) and “0” (not entailment). For our experiments we are only concerned with the entailment class.
LLM
We also experiment with using an LLM to do the citation recovery task. We zero-shot prompt an LLM to recover the sentences ($X_{pred}^{mask}$), using an instruction similar to what we provide to the human annotators (Section 3.4). For the input, we include the evidence passage ($e_{k}$) and all explanation sentences in the prompt. See Appendix Table 13 for an example prompt.
In terms of LLMs, we use the same set of models that we used for generating explanations (i.e. GPT-4, GPT-3.5 and LLaMA2), and further include several smaller LLMs: LLaMA3.1 (Dubey et al., 2024), Mistral (Jiang et al., 2023) and Gemma (Mesnard et al., 2024). [^9] Detailed LLM versions: LLaMA3.1 = Llama-3.1-8B-Instruct, Mistral = Mistral-7B-Instruct-v0.3, Gemma = gemma-1.1-7b-it. All LLMs have at least 8K context window, and as such they fit the full prompt containing instructions, evidence set, and explanation.
Evaluation Setting
We perform automatic annotation under two settings: (1) sampling: for each claim, only one evidence passage is selected for masking – this is the approach we used in the human annotation in Section 3; and (2) full: for each claim, we mask and evaluate all evidence passages, one at a time. The full setting means we are assessing the attribution accuracy for every sentence that contains a citation marker in the explanation, and providing a more complete picture of the quality of the generated explanations (noting that this wasn’t feasible with human evaluation). Note that in this setting, for each claim we have multiple set precision, recall and F1 scores (one for each evidence passage; Section 3.3), and we aggregate them by computing the mean.
Figure 4: The agreement between automatic annotators vs. human annotators. “Generator” = models that are also used as (explanation) generators; $\sim$7b LLM” refers to open-source LLMs with around 7B parameters; “NLI PLM” indicates PLMs used to conduct pairwise NLI, as described in Section 4.
4.1 Agreement with Human Annotation
To understand how reliable LLMs and NLI models are as automatic annotators, we first assess their agreement with human annotations. We compute the agreement between automated annotator and human annotator using krippendorff’s alpha in a similar manner to that described in Section 3.5. But here we compute the agreement between two sets of annotation results: (1) an automatic annotator’s results (e.g. GPT-4’s); and (2) aggregate results from all human annotators. [^10] That is, we take the union of all recovered sentences ($X_{pred}^{mask}$) for each human annotator into a larger set. We also standardise any labels outside the ground truth set ($X_{cit}^{mask}$) to a common label, as we are mostly interested in understanding the extent to which the automatic annotator selects the ground truth label when human annotators do so. [^11] For example, if the ground truth set ($X_{cit}^{mask}$) is {9}, and the predicted set ($X_{pred}^{mask}$) from the automatic and human (grouped) annotator is {7, 9} and {5, 9} respectively, we will post-process both predicted sets into {-2, 9}.
We present the agreement results in Figure 4. We can see LLM annotators have a much higher correlation with human annotators than NLI annotators. Due to the nature of the NLI input, the hypothesis is an explanation sentence that is taken out of context of the whole explanation, and may not always be interpretable. The LLM also assesses each explanation sentence individually, but they do this by taking into account the whole explanation context (as it is provided as part of the prompt). Given these results, we focus on using LLMs as annotators for the remainder of our experiments.
4.2 Automatic Evaluation Results
We now present citation recovery results using LLMs as annotators for sample in Table 5 and all in Table 6. For the former we also include the human annotator results for comparison.
Comparing Automatic Annotator vs. Human Annotator
In Table 5, we see that most automatic annotators have high agreement with human annotators: GPT-4 has the best attribution quality, followed by GPT-3.5 and LLaMA2. That said, if we look at the magnitude of the attribution quality, GPT-3.5 seems to overestimate this (i.e. its F1 numbers are much higher than those for humans), and this correlates with the agreement performance we saw in Section 4.1 where GPT-3.5 has a lower agreement. GPT-4 produces results that best align with human annnotators, and this is also reflected in Section 4.1 where GPT-4 has the highest agreement.
Ann Gen Evi Src GPT-4 GPT-3.5 LLaMA2
Attribution F1
GPT-4 Human 0.69 0.60 0.50
Machine 0.62 0.57 0.55
GPT-3.5 Human 0.83 0.73 0.61
Machine 0.75 0.69 0.60
LLaMA2 Human 0.65 0.60 0.47
Machine 0.57 0.51 0.47
LLaMA3.1 Human 0.57 0.49 0.45
Machine 0.56 0.46 0.40
Gemma Human 0.69 0.58 0.49
Machine 0.59 0.52 0.46
Mistral Human 0.58 0.49 0.42
Machine 0.49 0.51 0.44
Human Human 0.63 0.52 0.49
Machine 0.74 0.53 0.49
Table 5: Evidence attribution results for the sample setting. The “Evi Src” column indicates whether evidence is Human- or Machine-selected, “Ann” the automatic annotators and “Gen” the explanation generators.
Ann Gen Evi Src GPT-4 GPT-3.5 LLaMA2
Attribution F1
GPT-4 Human 0.67 0.61 0.50
Machine 0.60 0.50 0.40
GPT-3.5 Human 0.80 0.75 0.64
Machine 0.74 0.57 0.45
LLaMA2 Human 0.59 0.58 0.48
Machine 0.53 0.41 0.34
LLaMA3.1 Human 0.47 0.47 0.40
Machine 0.44 0.36 0.26
Gemma Human 0.67 0.63 0.50
Machine 0.55 0.47 0.34
Mistral Human 0.54 0.49 0.39
Machine 0.48 0.40 0.30
Table 6: Evidence attribution results for the full setting.
Comparing sample vs. full
Looking at both Table 5 and Table 6, we see a similar trend across both sample and full, where GPT-4 is consistently the best generator, followed by GPT-3.5 and LLaMA2. Interestingly though, if we look at the impact of evidence source (“human” vs. “machine”), for sample we see a small improvement from “machine” to “human”, but this gap widens further in full (for generator GPT-3.5 and LLaMA2 in particular), suggesting that human-selected evidence does produce better explanations. This observation contrasts what we saw previously in Section 3.5 (where machine-selected evidence is on-par with human-selected evidence), and shows that it is important to assess attribution for every sentence in the explanation to get the complete picture.
To provide a more intuitive understanding of the overall transparency of the generated explanations, we present Table 7 to show the proportion of explanations where every sentence attributes the source accurately (evaluated by GPT-4). We observe a large gap between “human” and “machine”, again reinforcing that human-curated evidence leads to more transparent explanations.
Discussion
Putting these results together, we can conclude as follows: (1) GPT-4 is the best model for generating fact-checking explanations; (2) there is still much room for improvement for explanation generation, because even with human-curated evidence, only 31% of GPT-4’s explanations attribute the sources accurately; (3) GPT-4 is the best automatic annotator for assessing evidence attribution; and (4) the selection of evidence is crucial, and we found that human-selected evidence tends to produce higher quality explanations across different LLM generators.
Ann Gen Evi Src GPT-4 GPT-3.5 LLaMA2
Transparency Proportion
GPT-4 Human 31% 23% 17%
Machine 9% 5% 4%
Table 7: The proportion of Explanation where all sentences are transparent. We use a threshold of 0.6 for attribution F1 score to classify instance as “transparent”.
5 Conclusion
In this paper, we explore evaluating evidence attribution for fact-checking explanations. We implement a novel protocol using both human annotation and automated methods, and show that our LLM-based annotations have stronger correlation with human annotations compared to NLI models. While the best-performing LLM shows promising results, there is still substantial room for improvement in explanation generation. Our findings highlight the importance of evidence selection, as human-curated evidence tends to produce more attributable explanations. Ultimately, our research provides insights into enhancing AI trustworthiness and paves the way for future developments of transparent explanation generation for fact-checking.
Limitations
This work only evaluates the evidence attribution of the generated explanations, i.e. how faithful they are based on the sources. We want to note that this evaluation does not assess factuality, which checks whether an explanation is factually correct, that said, there are already many studies on factuality, and our evaluation is orthogonal to these studies. Additionally, while we tried our best to craft the prompts, there remains the question of whether evidence attribution can be further improved with more prompt engineering. The PolitiHop dataset may contain out-of-date information, given it was developed a few years ago. As such, this may conflict with the LLM’s parametric knowledge for certain claims. We saw this in our analyses where LLMs ignore evidence passages and generate statements based on their intrinsic knowledge. Also, the claims and evidence passages in PolitiHop are generally short single sentences. In practice, claims and evidence passages are likely to be longer.
Ethics Statement
Appendix A Assessing Perceived Utility
Although not the main focus of this paper, in addition to Evidence Attribution, we also assess Explanation Utility as a complementary metric. Utility captures whether users find an explanation helpful in clarifying the claim Liu et al. (2023); Gao et al. (2023).
Utility Evaluation
Utility measures to what extent the generated explanation clarifies a claim. Though a five-point Likert scale is commonly used (Liu et al., 2023; Gao et al., 2023), Ethayarajh and Jurafsky (2022) found that averaging using the Likert scale can result in a biased estimate. As such we use Direct Assessment (Graham et al., 2013) where annotators rate on a (continuous) sliding scale from 00100100100100, where $100$ = best. We ask annotators the following question: How helpful is the explanation in clarifying the truthfulness of the claim?. We aggregate the judgements of multiple annotators for each explanation by computing the mean. We also applied a Bayesian model for utility score calibration, but it showed a similar tendency, so we decided to use the original scores.
Utility Score Calibration
@@ Appendix A Assessing Perceived Utility
Utility Results
High utility doesn’t necessarily imply high evidence attribution. Although we found a a general pattern between utility and evidence attribution (Table 9), GPT-3.5 achieved the highest utility score when using human-selected evidence (72.47), even though its attribution score is much lower compared to GPT-4 (0.52 vs. 0.63). After filtering for transparent explanation with a threshold of $F1=1.0$ and $F1\geq 0.9$, GPT-4 achieved the highest utility score, as shown in Table 10.This indicates that high utility GPT-3.5 explanations are not attributable. Overall, this demonstrates that utility and citation represent two distinct qualities and has an important implication: an explanation that appears helpful can actually still be misleading.
Appendix B Annotation Entropy
We used another metric Entropy (Shannon, 1948) to measure the randomness and the degree of uncertainty in the system:
$H=-\sum{{p_{k}}\;{log}\;{p_{k}}}$ (8)
Appendix C Prompts
We developed and optimized prompt for our task. We present evidence selection prompt in Table 14, explanation generation prompt in Table 12 and explanation evaluation prompt for LLM as annotators in Table 13.
Appendix D Experiment Details
We ran all offline models on 4 Nvidia A100 GPUs in a data parallel fashion. Explanation generation with GPT-4 and GPT-3.5 takes around 2 hours and LLaMA2 takes around 4 hours. The cost of generating explanations was \$49 (US Dollar) with GPT-4 and \$10 (US Dollar) with GPT-3.5. We present full evidence attribution results for sample setting in
@@ Appendix D Experiment Details
Table 14: Prompts for evidence selection.
Appendix E Scientific Artifacts
Appendix F Annotation Interface